# Classical origins of Landau-incompatible transitions

### v1

@@ B.5 $Q=2$ phase diagrams
Figure S4: Phase diagrams for the $Q=2$ Hamiltonian of  Eq. 20. Arrowheads indicate the pinned values of $\langle\theta\rangle$ First and second order transitions are indicated by solid and broken lines. A direct transition between Landau-incompatible phases requires tuning two parameters.
We now consider the $Q=2$ model of Eq. 20 whose phase diagrams are shown in Fig. S4. The symmetry group of Eq. 21 is abelian,
@@
The phase diagrams for various $\gamma_{2}$ values are shown in Fig. S4. We see that a direct Landau-incompatible transition is not stable, and needs fine-tuning two parameters for $\gamma_{2}\neq 0$, i.e., $\gamma_{1}=0$ and $T=T^{\gamma}$ or $\gamma_{1}=\gamma_{2}=0$. However, this fine-tuning can occur accidentally quite naturally in quantum models with aesthetic appeal. For instance, setting $\gamma_{\ell\geq 2}=0$ in Eq. 28 yields the XYZ model [3] with a staggered magnetic field,
@@
A final interesting feature of Fig. S4 is the presence of symmetry-enriched Ising criticality [15, 16]. Transitions between PM and p-SSB± phases belong to the Ising universality class. However, symmetries act on the two branches in different ways. This is seen from the fact that the order parameters that correspond to the primaries of the Ising CFT transform as different irreducible representations of the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry in Eq. 36. It is sufficient to track the primary field $\sigma$. For the transition to p-SSB+, this corresponds to the lattice operator $\sigma\sim\cos(\theta)$ which is charged under ${\mathcal{R}}$ but not ${\mathcal{T}}$, while for the transition to p-SSB-, this corresponds to $\sigma\sim\sin(\theta)$ which carries both ${\mathcal{R}}$ and ${\mathcal{T}}$ charges. The distinct charge assignments mean that the two Ising CFT branches cannot be connected trivially. In Fig. S4, they pass through a different universality class with $c=1$.
B.6 $Q=3$ phase diagrams
@@
Figure S5: Phase diagrams for the $Q=3$ Hamiltonian of  Eq. 20. Arrowheads indicate the pinned values of $\langle\theta\rangle$ First and second order transitions are indicated by solid and broken lines.
The phase diagrams for the $Q=3$ Hamiltonian are shown in Fig. S5. The two p-SSB± regions correspond to the same phase and the two branches of transitions between them and the PM belonging to the 3-state Potts universality class are same symmetry-enriched versions. As the unnecessary critical line terminates, the two branches can be connected smoothly.
B.7 $Q=4$ phase diagrams
@@
Figure S6: Phase diagrams for the $Q=4$ Hamiltonian of Eq. 20. Arrowheads indicate the pinned values of $\langle\theta\rangle$. First- and second-order transitions are indicated by solid and broken lines. The lines marked ‘AT’ have varying critical exponents and correspond to the orbifold branch of the conformal manifold that describes the critical Ashkin-Teller model.
@@
The two lines of transitions between PM and p-SSB± seem to have different symmetry enrichments seen by tracking the symmetry charges of the order parameters for SSB± $\cos(2\theta)$ and $\sin(2\theta)$. But, as explained in Ref.[15] and seen in Fig. S6, they are smoothly connected via the KT point without leaving the conformal manifold and should nominally be considered as belonging to the same symmetry-enriched class.
B.8 $Q\geq 5$ phase diagrams
(a) $\gamma<0$
(b) $\gamma>0$
Figure S7: Phase diagrams for the $Q\geq 5$ Hamiltonians of Eq. 20. First- and second-order transitions are indicated by solid and broken lines.
@@ B.8 $Q\geq 5$ phase diagrams
Appendix C Comments on anomalies

### v2

@@ I.5 $Q=2$ phase diagrams
Figure 9: Phase diagrams for the $Q=2$ Hamiltonian of  Eq. 19. Arrowheads indicate the pinned values of $\langle\theta\rangle$ First and second order transitions are indicated by solid and broken lines. A direct second order transition between Landau-incompatible phases requires tuning two parameters.
We now consider the $Q=2$ model of Eq. 19 whose phase diagrams are shown in Fig. 9. The symmetry group of Eq. 20 is abelian,
@@
The phase diagrams for various $\gamma_{2}$ values are shown in Fig. 9. We see that a direct Landau-incompatible transition is not stable, and needs fine-tuning two parameters for $\gamma_{2}\neq 0$, i.e., $\gamma_{1}=0$ and $T=T^{\gamma}$ or $\gamma_{1}=\gamma_{2}=0$. However, this fine-tuning can occur accidentally quite naturally in quantum models with aesthetic appeal. For instance, setting $\gamma_{\ell\geq 2}=0$ in Eq. 28 yields the XYZ model [3] with a staggered magnetic field,
@@
A final interesting feature of Fig. 9 is the presence of symmetry-enriched Ising criticality [19, 20]. Transitions between PM and p-SSB± phases belong to the Ising universality class. However, symmetries act on the two branches in different ways. This is seen from the fact that the order parameters that correspond to the primaries of the Ising CFT transform as different irreducible representations of the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry in Eq. 36. It is sufficient to track the primary field $\sigma$. For the transition to p-SSB+, this corresponds to the lattice operator $\sigma\sim\cos(\theta)$ which is charged under ${\mathcal{R}}$ but not ${\mathcal{T}}$, while for the transition to p-SSB-, this corresponds to $\sigma\sim\sin(\theta)$ which carries both ${\mathcal{R}}$ and ${\mathcal{T}}$ charges. The distinct charge assignments mean that the two Ising CFT branches cannot be connected trivially. In Fig. 9, they pass through a different universality class with $c=1$.
I.6 $Q=3$ phase diagrams
@@
Figure 10: Phase diagrams for the $Q=3$ Hamiltonian of  Eq. 19. Arrowheads indicate the pinned values of $\langle\theta\rangle$ First and second order transitions are indicated by solid and broken lines.
The phase diagrams for the $Q=3$ Hamiltonian are shown in Fig. 10. The two p-SSB± regions correspond to the same phase and the two branches of transitions between them and the PM belonging to the 3-state Potts universality class.
I.7 $Q=4$ phase diagrams
@@
Figure 11: Phase diagrams for the $Q=4$ Hamiltonian of Eq. 19. Arrowheads indicate the pinned values of $\langle\theta\rangle$. First- and second-order transitions are indicated by solid and broken lines. The lines marked ‘AT’ have varying critical exponents and correspond to the orbifold branch of the conformal manifold that describes the critical Ashkin-Teller model.
@@
The two lines of transitions between PM and p-SSB± seem to have different symmetry enrichments seen by tracking the symmetry charges of the order parameters for SSB±, $\cos(2\theta)$ and $\sin(2\theta)$. But, as explained in Ref. [19] and seen in Fig. 11, they are smoothly connected via the KT point without leaving the conformal manifold and should nominally be considered as belonging to the same symmetry-enriched class.
I.8 $Q\geq 5$ phase diagrams
(a) $\gamma_{1}<0$
(b) $\gamma_{1}>0$
Figure 12: Phase diagrams for the $Q\geq 5$ Hamiltonians of Eq. 19. First- and second-order transitions are indicated by solid and broken lines.
@@ I.8 $Q\geq 5$ phase diagrams
Appendix II Comments on anomalies
@@ II.1 Anomalies in quantum systems
Deconfined criticality is sometimes associated to the microscopic symmetries being realised anomalously [4] either on the ultraviolet [3, 41, 42, 43] or infrared [11, 12, 62] degrees of freedom. Note that this prejudice is not universally held, as evidenced by the search for a direct DQC transition between the Néel and valence-bond solid (VBS) phases on the honeycomb lattice [63, 12] which is not constrained by the LSM theorem. However, recent work [64] has indicated that this transition may in fact not exist. Our work proves that even if this particular DQC is not present in the honeycomb lattice, there is no fundamental obstruction to observing Landau-incompatible transitions on the honeycomb or other LSM-trivial lattices.
II.2 Anomalies in classical systems
A well known result is that the ground state phases of a $d$-dimensional quantum system can be reproduced by an appropriate $d$+1 dimensional classical system [65]. When the quantum system is anomalous, and its phase diagram does not contain a trivial, symmetric phase, how does this manifest in the corresponding classical system? In particular, at high enough temperatures, classical systems are expected to restore all symmetries and produce a trivial disordered phase, seemingly in contradiction to the anomaly constraint of the corresponding quantum system  [^5] We thank the anonymous referee for posing this interesting question! The resolution of this paradox comes from the observation that the classical system corresponding to an anomalous quantum system can be constrained. For example, consider the XXZ spin chain which has a well-known LSM anomaly [12]. Its corresponding classical system is the six-vertex model [66, 67]. If we consider the partition function for this model with unit Boltzmann weights, representing the infinite temperature limit, we have
${\mathcal{Z}}_{6v}=\sum_{\mathcal{C}_{6v}}1.$ (38)
This is in a disordered phase characterised by pseudo long-range order and algebraic correlations (it corresponds to the eight-vertex model at a critical temperature) [66], and hence does not correspond to the trivial disordered phase which has exponentially decaying correlations. The six-vertex constrained nature of the configuration space $\mathcal{C}_{6v}$ over which the partition function sum is performed introduces non-trivial correlations even when all Boltzmann weights are unity. This is also true of other constrained systems, for example, two-dimensional dimer models [68, 69]. On the other hand, if the configuration space were to be constraint-free, as is true of the classical Ising and XY model, the partition function with unit Boltzmann weights would indeed be trivial with exponentially decaying correlations.